A Bibliography of Publications in Java Report: The Source for Java Development

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 November 2010 Version 1.10

Abstract

This bibliography records publications in the Java Report magazine.

Title word cross-reference

1.0 [Bru97]. 1.1 [Gea97a]. 100 [Gil98b]. 1997 [Wat97b].

2nd [Jor98a].

’97 [Wat97c].

Abridged [Ano96a]. abuse [BC98c]. Acceptance [Bro98d]. Access [BR96]. Across [Jol98]. Activate [Ens98].

ActiveX [Til97a]. Adaptors [Eas97]. add [Won97c]. add-on [Won97c]. Adding [Fei98d]. Address [Sei97]. advantages [DB97a]. Advertiser [Ano96b]. Advice [Ano97j]. Affect [Sco96a]. Affordable [Tho97f]. Agent [Far97]. Aiken [Tho97e]. Aims [Gil97f]. Alive [Wat97a]. All-Java [Gil97f]. alphaBeans [Ano98a]. Alternatives [Bro98e]. Analysis [Tro96]. Annual [Jor98a]. Answer [Oak97h]. Answering [Sco96b]. Answers [Ano98a]. Antidotes [Bow97a]. App [Gil98c]. Applet [Ano97q]. Applets [MF97]. Application [Jor97b]. Applications [BR96]. Applying [Gam96]. Approach [WN96]. Apps [Gil98b]. Architecture [Far97]. areas [Mee96]. aren’t [Mee96]. Art [Swa97]. asked [Gea97b]. Ass [Sch97]. Assembly [Ses98d]. Assembly-Line [Ses98d]. Assessing [Sco97a]. Assignment [Jol98]. Atomic [Jol98]. Authenticode [Fly97]. Avenue [Til96b]. Avoiding [DB97b]. Aware [Bro98f]. AWT [Gea97b].

Back [Oak98a]. Bar [Bro98a]. based [Oak97a]. Basic [Adl98]. Basics [Oak98a]. Battle [Gil98d]. Beachhead [Gil97i]. Bean [Bro98a]. Bean-Aware [Bro98f]. Beans [Mis97, Hou98a]. bedfellows [Til96b]. behavior [Oak97b]. Behind [Mat97a]. being [SR97]. best [Sco97b]. Bets [Gil97p]. Better [MP98]. between [Til96b]. billing [Til96b]. Birthday [Deu97g]. Bonus [Ano97h]. Book [Bar96]. boost [Mat97a]. Borland [Gil97h]. Brews [Gil98c]. Browser [Gil98d]. buffering [Mee96]. Build [Fei98d]. Builder [Fei97]. Building [Tho97a]. Bulletproof [NST97]. Burden [Oak98b]. Business [BR96]. Buys [Gil98b]. Buzz [Ano96f]. byte [Oak97j].

C [Joy96, Mar97]. C/C [Joy96]. CAB [Zho97]. Café [Bru97, Fit96]. CAGE [Gra97h]. can [Sco97c]. capture [Jov96]. Cards [Deu97a]. Cartoons [Mey96]. Case [Gil97d]. Cash [Gil98a]. Cautionary [Dic98]. Certification [Dou97]. Certified [Ano97p]. Chains [Smi98]. Changing [Oak97b]. CICS [Gil98b]. cities [Wer97a]. Class [Tro96]. Classified [Ano96a]. Client [Sco96c]. Client-Side [Sco96c]. closer [Mey96]. Coda [Gra97h]. code [Oak97j]. Coding [Kal98]. Coffee [Ano96d]. Come [Gil98b]. Coming [Gil97g]. command [Til96b].

Commandexecutor [Ano98b].

Commitment [Wat98b]. communication [Jov96]. Comparison [Mar97]. competitive [BR96]. Compiler [Gil98d]. complex [BR96]. Compliance [Gil98b]. Component [Tho98a]. Component-Based [Tho98a]. Components [Bow97b]. Computers [Meh97]. computing [Deu97d]. concept [Fra98]. Conceptual [Lar98a]. Conference [Var96, Var96]. Connection [Wer97b]. Connectivity [Jep97]. considerations [Oak96a]. Considered [Alp98]. Construct [Tho97a]. Constructing [Hun97]. Consumers [Won97h]. Control [Boy97]. CORBA [Ses98c]. Corel [Gil97f]. Corp [Gra97h]. Corporate [McC96]. cost [Jov96]. Crafting [Cel97a]. Crawls [Ano98c]. Creating [Sco96c]. Critical [Mar97]. Cubed [Gil98c]. Cure [MP98]. currently [Mee96]. Customer [BR96].

Data [Jol98]. Database [Jov96].

Database-Driven [Jov96].

databases [DB97c]. degree [Jor98b]. Delegation [Sev97]. Deliver [Ano98c]. Delivered [Won96b]. Delivers [Pep97]. Delivery [Mil98]. Deployment [Zin98a]. Derby [Jor98a]. design [Tro96]. Designer [Gra97h]. Desk [Ses98c]. Desktop [Kal98]. Developer [Wat97b]. Developers [Var96]. Developing [Jov96]. Development [Jav96]. Devices [Gil98b]. Digitivity [Gra97h]. Directly [Gea96]. Director [Ano97k]. Directory [MH98b]. Distributed [BR96]. do [Sco96b]. Dolby [Wat97a]. Domino [Wat97b]. double [Mee96]. Dreams [Gil98d]. Driven [Jov96]. dynamic [BR96].

Easier [Cow97]. EC [Gil98c]. ecWorkRouter [Gil98c]. edges [Til97c]. Edition [Tho97g]. Editor [Ano97m, Ano96e]. Editorial [Maj97]. Editors [Ano97p]. EDS [Gil98b]. Education [Opd97]. Efforts [Ano98a]. EJB [Gil98d]. Embedding [Gea96]. Embodies [Bro98c]. emerging [Jov96]. end [BC98a]. Enemy [Kal98]. Engine [Ano98c]. Enhances [Ano98c]. Enter [Sco96c]. Enterprise [Kap97a, Gra97h]. Enumerating [Oak97d]. environments [BR96]. Espeset [Sch97]. Evaluating [ALW97]. evangelist [Ano96d]. event [Gea97e]. events [Oak96a]. Every [Oak97h]. Exam [HR97]. Exception [Won97c]. Execution [Kuz97]. Exercising [Dis98a]. Experiment [Mey96]. Exploring [Jov96].

Facilitating [Kap97a]. Facility [Fei98d]. factory [Oak97b]. falling [Til97b]. Faster [Bro98b]. Fastest [Gil98d]. Fatherly [Ano97j]. Fear [Mor97]. feature [Won97c]. features [Won97g]. Feigenbaum [Fei98c]. few [Oak96a]. Fighting [Ses98b]. Filling [Fei97]. Financial [Gil97d]. Fire [Ses98b]. First [Kap97b]. Flexible [MH98b]. Focus [Gil97p]. follow [Til98a]. Fooling [Mey96]. Framework [Cel98]. Free [Ano98a]. French [Mey96]. frequently [Gea97b]. Fujitsu [Gil98c]. functions [BR96]. Future [Ses98d].

GemStone [Gil97p]. Generation [Opd97]. Generic [Kap98b]. Get [Sut98]. Gets [Gil98c]. Getting [Tro96]. Giant [Gil97f]. Goes [Jae97]. Got [Ano98a]. Graphic [Gea96]. graphical [Mee96]. Great [Bro98d]. Guest [Deu97g]. guidelines [Til96b]. GUIs [Cel97a].

Handlers [Mal98]. handling [Oak96a]. Harmful [Alp98]. HashMap [Oak98b]. Hashtable [Oak98b]. Hello [Mat97b]. Help [Bro98a]. Helping [BC98a]. Here [Gil98b]. higher [Jor98b]. Hole [Fei97]. Holes [Bow97a]. Hopes [Gil97f]. Hostile [MF97]. HP [Gil97g]. HTTP [Juc97].

IBI [Gil98c]. IBM [Ano98c]. IDEs [Zin98b]. IFC [Cel97b]. image [Won97a]. images [Mee96]. Immunizes [Gil98b]. Imperial [Tho98a]. importance [Bow97c]. improve [Jov96]. Improvements [Til96a]. Improving [Mee96]. Increasing [BR96]. Incremental [Mis97]. independent [Fra98]. Index [Ano96b]. Industry [Gil97d]. Information [Ano98c]. InnerClasses [Ano97p]. Innovative [Tho97f]. Inprise [Wat98b]. Insider [Mat97b]. instanceof [BC98c]. Integrated [ALW97]. Integrates [Kal98]. Integrating [Wat98b]. Integration [Gil98c]. Intelligence [McC96]. interactive [Mey96]. Interface [WN96, Jor97a]. Interfaces [Hun97]. International [Jae97, Var96]. Internet [Jov96]. interview [Ano96d]. Introduction [Jor97a]. Invocation [Oak97c].

J [Fly96, Wer97b]. J/SQL [Wer97b]. Java [Var96, Sch97, Jav96]. Java-Based [Gil98c]. JavaBean [MS97]. JavaBeans [Eas97]. JavaOS [Gil98d]. JavaScript [Tho97e, Sco96c]. JavaSpaces [Sha97b]. jBusiness [Zin98a]. JClass [Ano97q]. JDBC [Wer97b, Tho97d]. JDBTools [Gra97h]. JDesigner [NST97]. JDK [Won97g]. JFC [Gea97c]. Jini [Gil98d]. JNDI [MH98b]. Jobs [Man96]. Junkie [Ses98e]. just [Won97c].

Key [Gil97h]. Kick [Sch97]. King [Ses98c]. Know [Fly97].

Languages [Fei98d]. Latest [Bro98c]. layers [Sco97a]. layout [Won96a]. Lead [Gil98b]. Learning [Tho97f]. Legacy [Gil98c]. Letters [Ano97m]. Life [Won97f]. Lightweight [Gil98d]. Line [Ses98d]. LiveTable [Ano97q]. Lobby [Sut98]. Logic [Zin98a]. look [Won96a]. Looking [Won97g]. Lotus [Wat97b]. Lure [Jep97].

Machine [Cra98]. Macintosh [Fit96]. Made [McM98]. Madhu [Sut98]. Madison [Til96b]. Magic [Mil98]. mail [Oak97a]. Mainframes [Deu97a]. Make [Che97]. makes [Gea97a]. Making [O’B97]. Management [Gil98a]. manager [Won96a]. managers [Won96a]. Marketplace [Tho97c]. marquees [Mee96]. Marriage [Sco96d]. Matsumura [Ano96d]. May [Ano98a]. Means [Sco96b]. media [Won97d]. Memories [McM98]. MerzCom [Gra97h]. MerzScope [Gra97h]. meta [Fra98]. meta-model [Fra98]. Method [Oak97c]. Methods [Won97i]. Microsoft [Fly96]. Mid [Gil97p]. Mid-Tier [Gil97p]. Miko [Ano96d]. misused [SR97]. Mobile [Far97]. Model [WN96]. Model-View [WN96]. Model/View/Controller [Gea97d]. Modular [Hun97]. Money [Til96b]. Moving [Won97h]. Multi [BC98b]. Multi-Threaded [BC98b]. Multimedia [Mey96]. Multithreaded [DH98]. multithreading [Gea98].

Naming [MH98b]. Natural [McC96]. Nears [Gil98d]. Need [Fly97]. net [Til97b]. NetDynamics [Gil98b]. Netiva [Pla97]. NetRexx [Cow97]. Netscape [Cel97b]. Network [Meh97]. Networking [Maj97].

News [Gra97a]. Next [Opd97]. No [Oak97h]. Non [Gil98b]. Non-Compliance [Gil98b]. Notable [Tho97c]. Note [Ano96e]. notify [Won97i]. Novera [Zin98a]. Numbers [Dic98].

Object [Tro96]. Object-Oriented [Tro96]. Objects [Tro96]. Objectshare [DCMR96]. Observer [Smi98]. observers [Won97a]. ODMG [Jor97a]. Offers [Gil98d]. Office [Gil97f]. Once [Kap98a]. One [Til96b]. Online [Mey96]. onto [Mee96]. Open [MH98b]. Optimizing [Kuz97]. option [Dis98a]. Oracle [Esk96]. Oriented [Tro96]. OTM [Gil98b]. OTMs [Gil98b]. Overtake [Gil97f]. own [Won96a].

packages [Oak97e]. Page [Mil98]. pain [Wer97b]. ParcPlace [DCMR96]. Parity [Gil98d]. Part [Gea97c]. PARTS [Zin98b]. pattern [DB97a]. Patterns [Gam96]. Controller [Gea97d]. View [Gea97d]. performance [Mat97a]. Perks [Ano98a]. Persistence [How97, Gil98c]. Persistent [O’B97]. Peter [Tho97e]. Pitfalls [Won97i]. Pizzazz [Cel98]. PJava [Gil98d]. Plan [Tho97a]. Platform [Gil98c]. Plug [Bro98c]. Plug-in [Bro98c]. Political [Sut98]. Port [Boy97]. portability [Jor98b]. Post [Ano98a]. Poster [Ano97h]. Power [Dis97]. PowerJ [Nat97]. practical [Oak96a]. Pragmatic [Won96a]. praise [Oak97f]. premium [Til96b]. Preparing [HR97]. Prevention [MP98]. Primitive [Alp98]. Privileged [Som98]. Pro [Ano97q]. Processes [Lar98b]. Producers [Won97h]. Product [Fit96]. Productivity [Ano98c]. Products [Tho97c]. Professional [Zin98b]. Profile [McC96]. Program [Wer97c].

Programmer [Won96a]. Programmers [Joy96]. Programming [Bar96]. Promise [Bro98d]. Protecting [Oak97j]. Protocol [Mal98]. Proxies [Juc97]. proxy [DB97a]. Publishing [Wat97c]. Pull [Ano97h]. Pure [Gil98b]. Push [Gil98b]. Putting [Hou98a].

question [Sco96b]. questions [Gea97b]. Quirky [Zin98b]. Quiz [Bro98c].

RandomNoise [Gra97h]. real [Sco97c]. really [Oak97g]. Recruitment [Ano96g]. reduce [Jov96]. Refinements [Tro96]. Reflection [Bro98e]. Related [Fei98d]. relational [DB97c]. Release [Gil98d]. Reliable [Che97]. Rendering [Mee96]. Report [Jav96]. Representing [DB97c]. Rescuing [BC98c]. Resolution [Til98b]. Resource [Ano98c]. Resurgence [Gil97h]. revenue [Jov96]. Review [Fit96]. Revolution [Gil97e]. Rick [Sut98]. Right [Dic98]. RMI [Ses97]. Robertson [Wat97a]. Robust [Tho98a]. Rogue [Gra97h]. RogueWave [Pep97]. Ross [Sut98]. Run [Kap98a].

s [Dis98a]. S390 [Deu97a]. Safer [Bro98b]. Sanga [Gil97i]. Scenes [Cra98]. Scott [Oak96a]. SDK [Gil98b]. Search [Ano98c]. Secrets [Ree98]. secure [Oak97g]. Security [MF97]. Seeks [Gil97i]. Separate [Ano98a]. Separation [Bos98]. Serial [Boy97]. Serialization [Bro98e]. Served [Ses98a]. Server [Mal97]. Servers [Ano98a]. services [Til96b]. Set [Gil98d]. Seybold [Wat97c]. Ships [Gil98c]. Shockwave [Til97a]. should [Til96b]. Shouldn’t [Mor97]. Siddalingaiah [Sut98]. Side [Sco96c]. Signed [Som98]. Simple [How97]. Simpler [Bro98b]. Simplifies [Zin98a]. Simplify [Wat98b].

Simplifying [Jor97b]. Singleton [DB97d]. Site [Kap97b]. Sites [Tho97a]. Situation [Oak97h]. Slow [Bro98d]. Smart [Deu97a]. SMTP [Oak97a]. SMTP-based [Oak97a]. Soaked [Mos97]. software [Sco96b]. Solution [Til98b]. Solutions [Oak96a]. Some [Ano97j]. Sound [Wat97a]. Source [Jav96]. Space [Gil98b]. span [Won97f]. Speaks [Esk96]. Special [ALW97]. specialty [Til96b]. SQL [Wer97b]. Standards [Deu97f]. Started [Tro96]. State [Swa97]. statements [DB97b]. Status [Bro98a]. Status-Bar [Bro98a]. Step [Kap97b]. StoreTable [How97]. Story [Bro98c]. strange [Til96b]. strategies [Jov96]. strides [Gea97a]. Strong [Gil97g]. Studio [Bro98c]. Study [Gil97d]. Success [Bro98c]. Suite [Gil97f]. Sun [Ano96d]. SuperCede [Tho97g]. support [BR96]. Surprises [BC98b]. Survey [Ses98e]. Swing [Gea97c]. Symantec [Fit96]. Synchronization [Lew98]. System [Far97]. Systems [Gil98a].

Tailored [Fei98d]. Take [Gil98b]. Taking [Wer97b]. tale [Wer97a]. Talk [Ano96d]. Taming [Hou98b]. Tapping [Dis97]. Targeted [Ano98c]. Targets [Gil98b]. techniques [Oak96a]. Technology [Opd97, Tho98a]. Technology-independent [Fra98]. Test [Hou98a]. Thomas [Wat97a]. Thread [Bos98]. Threaded [BC98b]. threads [Oak97i]. Three [Ree98]. Three-Tier [Ree98]. Three-Tiered [Gil98a]. Throughout [Zin98b]. tidy [BC98a]. Tier [Gil97p]. Tiered [Gil98a]. Tips [Sha97a]. Tonny [Sch97]. tool [Sco97b]. Tools [Pep97]. trackers [Won97d]. Training [Mos97]. Trenches [Mey97]. True [Mey96].

Turning [MS97]. Tutor [Mee96]. TV [Gra97h, Mey96]. Two [Oak96a]. types [Oak97d].

ugly [Til97c]. Ultimate [Ses98e]. Understanding [Kal98]. Unique [Zin98b]. Upgrading [Kap97b]. URLConnection [Oak96b]. URLStreamHandlers [Mon97]. use [Sco96b]. User [Hun97]. Using [BR96].

value [Sco97c]. VBScript} [Tho97e]. very [Won96a]. Vibe [Gra97h]. View [WN96]. Views [Meh97]. Virtual [Sco96c]. VirtuFlex [Gra97h, Gra97h]. Visaj [Tho98b, Tho98a]. visible [Mee96]. Visient [Gil98d]. Visix [Gra97h]. Visual [Bru97, Fly96]. Visualization [Sco97c]. VRML [Sco96c]. vs [Boo97].

wait [Won97i]. Watch [Ano97i]. Wave [Gra97h]. way [Til98a]. Web [Wat97b, Jov96]. Website [Bar96]. Where [Kap98a]. Will [Sco96a]. within [Oak97d]. Without [Mos97]. work [Oak97e]. Workflow [Gil98c]. World [Sco96c]. Worlds [Tho97a]. write [Won96a]. Writing [Bro98f].

Y2K [Gil98b]. Year [Gil97e].

References

Adler:1998:URJ

[Adl98] Dan Adler. Using RMI and JDBC from Visual Basic. Java Report: The Source for Java Development, 3(??):??, November 1998. CO-DEN JREPFI. ISSN 1086-4660.

Alpert:1998:PTC

[Alp98] Sherman R. Alpert. Primitive types considered harmful. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660.

Albelli:1997:SRE

[ALW97] Gary Albelli, Eric Lazarus, and Steve Wirts. Special report: Evaluating integrated development environments. Java Report: The Source for Java Development, 2 (11):33–??, December 1997. CO-DEN JREPFI. ISSN 1086-4660.

AndersenConsulting:1997:CP

[And97] Andersen Consulting. Corporate profile. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

                                                  //www.sigs.com/publications/
                                                  docs/java/9701/java9701.toc.
                                                  html.

Anonymous:1996:AJC

[Ano96a] Anonymous. An abridged Java classified. Java Report: The Source for Java Development, 1 (6):??, November/December 1996. CODEN JREPFI. ISSN 1086-

                                                  4660. URL http://www.sigs.
                                                  com/publications/docs/java/
                                                  9611/java9611.toc.html.

Anonymous:1996:AIa

[Ano96b] Anonymous. Advertiser index. Java Report: The Source for Java Development, 1(3):64, July and August 1996. CODEN JREPFI.

ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Anonymous:1996:AIb

[Ano96c] Anonymous. Advertiser index. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9611/java9611.toc.html.

Anonymous:1996:CTI

[Ano96d] Anonymous. Coffee talk: An interview with Sun’s Java evangelist Miko Matsumura. Java Report: The Source for Java Development, 1(3):14, 16–18, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Anonymous:1996:EN

[Ano96e] Anonymous. Editor’s note. Java Report: The Source for Java Development, 1(3):4, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Anonymous:1996:JB

[Ano96f] Anonymous. Java buzz. Java Report: The Source for Java Development, 1(3):6–8, 10, 12, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/

                                        docs/java/9607/java9607.toc.
                                        html.

Anonymous:1996:R

[Ano96g] Anonymous. Recruitment. Java Report: The Source for Java Development, 1(3):77, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

                                        //www.sigs.com/publications/
                                        docs/java/9607/java9607.toc.
                                        html.

Anonymous:1997:AIa

[Ano97a] Anonymous. Advertiser index. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

                                        //www.sigs.com/publications/
                                        docs/java/9701/java9701.toc.
                                        html.

Anonymous:1997:AIb

[Ano97b] Anonymous. Advertiser index. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

                                        sigs.com/publications/docs/
                                        java/9702/java9702.toc.html.

Anonymous:1997:AIc

[Ano97c] Anonymous. Advertiser index. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

                                        sigs.com/publications/docs/
                                        java/9703/java9703.toc.html.

Anonymous:1997:AId

[Ano97d] Anonymous. Advertiser index. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Anonymous:1997:AIe

[Ano97e] Anonymous. Advertiser index.

Java Report: The Source for Java Development, 2(5):64–??, May 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9705/9705.toc.html.

Anonymous:1997:AIf

[Ano97f] Anonymous. Advertiser index.

Java Report: The Source for Java Development, 2(6):56–??, June 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9706/java9706.toc.html.

Anonymous:1997:AIg

[Ano97g] Anonymous. Advertiser index.

Java Report: The Source for Java Development, 2(8):64–??, September 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9709/java9709.toc.html.

Anonymous:1997:BPP

[Ano97h] Anonymous. Bonus pull out poster. Java Report: The Source for Java Development, 2(4):??, April 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Anonymous:1997:BW

[Ano97i] Anonymous. Book watch. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9703/java9703.toc.html.

Anonymous:1997:CTS

[Ano97j] Anonymous. Coffee talk: Some fatherly advice. Java Report: The Source for Java Development, 2(4):??, April 1997. CO-DEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9704/9704.toc.html.  Inter-

view with Java creator James Gosling, lead engineer for the Java/HotJava Systems.

Anonymous:1997:DNJ

[Ano97k] Anonymous. Director’s note: Java buzz. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

                                        //www.sigs.com/publications/
                                        docs/java/9701/java9701.toc.
                                        html.

Anonymous:1997:LEb

[Ano97l] Anonymous. Letters to editor. Java Report: The Source for Java Development, 2(5):6–??, May 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

                                        sigs.com/publications/docs/
                                        java/9705/9705.toc.html.

Anonymous:1997:LEa

[Ano97m] Anonymous. Letters to the editor. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9702/java9702.toc.
html.

Anonymous:1997:LEc

[Ano97n] Anonymous. Letters to the editor.

Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9711/java9711.toc.html.

Anonymous:1997:LEd

[Ano97o] Anonymous. Letters to the editor.

Java Report: The Source for Java Development, 2(11):6–??, December 1997. CODEN JREPFI. ISSN 1086-4660.

Anonymous:1997:LEJ

[Ano97p] Anonymous. Letters to the Editors: Java certified; InnerClasses. Java Report: The Source for Java Development, 2(8):??, September 1997. CODEN JREPFI. ISSN 1086-4660.

Anonymous:1997:PRJ

[Ano97q] Anonymous. Product review:

JClass LiveTable applet: and LiveTable Pro. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.

com/publications/docs/java/
9701/java9701.toc.html.

Anonymous:1998:ASG

[Ano98a] Anonymous. Application servers — got questions? we have some answers; Sun may separate Java standards efforts; IBM post free “alphabeans”; Java perks up software industry; and more. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/from_pages/ftp_javabuzz.
shtml.

Anonymous:1998:C

[Ano98b] Anonymous. Commandexecutor. Java Report: The Source for Java Development, 3(1):??, January 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9801/
commandexecutor.shtml.

Anonymous:1998:INJ

[Ano98c] Anonymous. IBM’s new Java resource enhances developer productivity search engine crawls the Web to deliver targeted Java information. Java Report: The Source for Java Development, 3(6):??, June 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9805/
jcentral.shtml.

Barrett:1996:BRW

[Bar96] Terry Barrett. Book review: Website programming with Java. Java Report: The Source for Java Development, 1(3):25–26, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Ball:1998:HOT

[BC98a] Steve Ball and John Miller Crawford. Helping objects to a tidy end. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9811/
html/from_pages/ftp_col2.shtml.

Ball:1998:MTA

[BC98b] Steve Ball and John Miller Crawford. Multi-threaded assignment surprises. Java Report: The Source for Java Development, 3(??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9809/
html/from_pages/ftp_col1.shtml.

Ball:1998:RIA

[BC98c] Steve Ball and John Miller Crawford. Rescuing instanceof from abuse. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9810/
html/from_pages/ftp_col1.shtml.

Boone:1997:CVJ

[Boo97] Barry Boone. C++ vs. Java software development. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9704/9704.toc.html.

Bosch:1998:TS

[Bos98] Peter Bosch. Thread separation. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660.

Bowman:1997:BRJ

[Bow97a] Charles F. Bowman. Book review: Java Security: Hostile Applets, Holes, and Antidotes. Java Report: The Source for Java Development, 2(5):29–??, May 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9705/9705.toc.html.

Bowman:1997:BOJb

[Bow97b] Charles F. Bowman. Business objects in Java: Business objects and components. Java Report: The Source for Java Development, 2(6):68–??, June 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9706/java9706.toc.html.

Bowman:1997:BOJa

[Bow97c] Charles F. Bowman. Business objects in Java: The importance of business objects. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Boydston:1997:SPC

[Boy97] David H. Boydston. Serial port control from Java. Java Report: The Source for Java Development, 2(6):53–??, June 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/java9706.toc.html.

Bandarpalle:1996:DBA

[BR96] Radha Bandarpalle and Raj Ratnakar. Distributed business applications using Java: Increasing customer access: How to support corporate business functions in dynamic, complex, and competitive environments. Java Report: The Source for Java Development, 1(3):46–50, 64, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Brookshier:1998:BSB

[Bro98a] Daniel Brookshier. Building a status-bar help bean. Java Report: The Source for Java Development, 3(3):??, March 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9803/
brookshier.shtml.

Brookshier:1998:EHS

[Bro98b] Daniel Brookshier. Exception handling: Simpler, faster, safer.

Java Report: The Source for Java Development, 3(2):??, February 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9802/
brookshier.shtml.

Brookshier:1998:JSS

[Bro98c] Daniel Brookshier. Java success story — Quiz Studio embodies the latest JDK, JFC components, and Java plug-in. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9811/
html/from_pages/ftp_realworld.
shtml.

Brookshier:1998:JGP

[Bro98d] Daniel Brookshier. JavaBeans — great promise, slow acceptance. Java Report: The Source for Java Development, 3(4):??, April 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9804/
brookshier.shtml.

Brookshier:1998:SRA

[Bro98e] Daniel Brookshier. Serialization and reflection — alternatives to using CORBA. Java Report: The Source for Java Development, 3(6):??, June 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9805/
brookshier.shtml.

Brookshier:1998:WBA

[Bro98f] Daniel Brookshier. Writing bean-aware applications. Java Report: The Source for Java Development, 3(3):??, March 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9803/
brookshier2.shtml.

Brugge:1997:PRS

[Bru97] John Brugge. Product review: Symantec’s Visual Café 1.0. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9703/java9703.toc.
html;   http://www.sigs.com/
publications/docs/java/9703/
9703.brugge.html.

Celestino:1997:CGU

[Cel97a] R. J. Celestino. Crafting GUIs using Netscape’s IFC. Java Report: The Source for Java Development, 2(6):39–??, June 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/9706.celestino.html;
http://www.sigs.com/publications/
docs/java/9706/java9706.toc.
html.

Celestino:1997:WNI

[Cel97b] R. J. Celestino. What Netscape’s IFC can do for you. Java Report: The Source for Java Development, 2(4):??, April 1997.

CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Celestino:1998:JMF

[Cel98] R. J. Celestino. The Java multimedia framework: Multimedia pizzazz. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660.

Cherlin:1997:HMJ

[Che97] Edward Cherlin. How to make Java more reliable. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9703/java9703.toc.html.

Cowlishaw:1997:NEJ

[Cow97] Mike Cowlishaw. NetRexx — easier Java programming. Java Report: The Source for Java Development, 2(6):45–48, June 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/java9706.toc.html.

Crawford:1998:BSJ

[Cra98]  Ronald Crawford, II.  Be-

hind the scenes of the Java 1.1 Virtual Machine. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9811/
html/from_pages/ftp_feature.
shtml.

Deugo:1997:PPAa

[DB97a] Dwight Deugo and Alan Benson. Practical patterns: The advantages of using the proxy pattern. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9702/java9702.toc.
html.

Deugo:1997:PPAb

[DB97b] Dwight Deugo and Allen Benson.

Practical patterns: Avoiding case statements. Java Report: The Source for Java Development, 2 (8):??, September 1997. CODEN JREPFI. ISSN 1086-4660.

Deugo:1997:PPR

[DB97c] Dwight Deugo and Allen Benson.

Practical patterns: Representing objects in relational databases. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Deugo:1997:PPS

[DB97d] Dwight Deugo and Allen Benson. Practical patterns: Singleton. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9701/java9701.toc.
html.

Dym:1996:CTO

[DCMR96] Richard Dym, Eric Clayberg, Tom Murphy, and Dan Rubel. Coffee talk: Objectshare and ParcPlace. Java Report: The Source for Java Development, 1 (6):??, November/December 1996. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9611/java9611.toc.html.

Deugo:1997:CTS

[Deu97a] Dwight Deugo. Coffee talk: From S390 mainframes to smart cards. Java Report: The Source for Java Development, 2(11):25– ??, December 1997. CODEN JREPFI. ISSN 1086-4660.

Deugo:1997:EN

[Deu97b] Dwight Deugo. Editor’s note. Java Report: The Source for Java Development, 2(6):4–??, June 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

                                        sigs.com/publications/docs/
                                        java/9706/java9706.toc.html.

Deugo:1997:ENB

[Deu97c] Dwight Deugo. Editor’s note: Building applications. Java Report: The Source for Java Development, 2(8):4–??, September 1997. CODEN JREPFI. ISSN 1086-4660.

Deugo:1997:END

[Deu97d] Dwight Deugo. Editor’s note: Distributed computing. Java Report: The Source for Java Development, 2(10):??, November 1997.

CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9711/java9711.toc.html.

Deugo:1997:ENG

[Deu97e] Dwight Deugo. Editor’s note:

Getting an education. Java Report: The Source for Java Development, 2(7):??, July/August 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9707/9707.toc.html.

Deugo:1997:ENS

[Deu97f] Dwight Deugo. Editor’s note:

Standards. Java Report: The Source for Java Development, 2 (11):4, December 1997. CODEN JREPFI. ISSN 1086-4660.

Deugo:1997:GEN

[Deu97g] Dwight Deugo. Guest Editor’s note: It’s Java’s birthday. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Deugo:1998:SMS

[Deu98] Dwight Deugo. From S390 mainframes to smart cards. Java Report: The Source for Java Development, 3(1):??, January 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9801/
coffeetalk.shtml.

DeRusso:1998:MEH

[DH98] Joe DeRusso, III and Peter Haggar. Multithreaded exception handling in Java. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/from_pages/ftp_feature.
shtml.

Dickey:1998:GNR

[Dic98] Ken Dickey. Getting the numbers right, A cautionary tale. Java Report: The Source for Java Development, 3(2):45–46, 48, 50, 52, February 1998. CODEN JREPFI. ISSN 1086-4660. Discusses interval arithmetic in Java for handling instabilities in polynomial evaluation.

Disbrow:1997:TPJ

[Dis97] Steven W. Disbrow. Tapping the power of JavaScript. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9703/java9703.toc.
html;   http://www.sigs.com/
publications/docs/java/9703/
9703.disbrow.html.

Disbrow:1998:EYO

[Dis98a] Steven Disbrow. Exercising your <option>s. Java Report: The Source for Java Development, 3 (??):??, October 1998. CODEN JREPFI. ISSN 1086-4660.

Disbrow:1998:MDJ

[Dis98b] Steven W. Disbrow. More dynamic JavaScript. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660.

Douglas:1997:CJD

[Dou97] Patricia J. Douglas. Certification for Java developers. Java Report: The Source for Java Development, 2(7):??, July/August 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9707/9707.toc.html.

Eastham:1997:EAJ

[Eas97]  Daniel A. Eastham.  Event

adaptors in JavaBeans. Java Report: The Source for Java Development, 2(6):33–??, June 1997. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9706/9706.eastham.
html;   http://www.sigs.com/
publications/docs/java/9706/
java9706.toc.html.

Enseling:1998:AYB

[Ens98] Oliver Enseling. Activate your beans. Java Report: The Source for Java Development, 3(3):??, March 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9803/
enseling.shtml.

Eskow:1996:CPO

[Esk96] Dennis Eskow. Corporate profile:

Oracle speaks. Java Report: The Source for Java Development, 1 (6):??, November/December 1996. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9611/java9611.toc.html.

Ethington:1998:DOC

[Eth98]  Russ Ethington.  Distributed

object computing. Java Report: The Source for Java Development, 3(??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9809/
html/from_pages/ftp_feature.
shtml.

Farley:1997:MAS

[Far97] Steven R. Farley. Mobile agent system architecture. Java Report: The Source for Java Development, 2(5):39–??, May 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Feigenbaum:1997:FHJ

[Fei97] Barry Feigenbaum. Filling a hole in Java: A Java class builder approach. Java Report: The Source for Java Development, 2(11):71– ??, December 1997. CODEN JREPFI. ISSN 1086-4660.

Feigenbaum:1998:CJCa

[Fei98a] Barry Feigenbaum. Code from: Java class builder. Java Report: The Source for Java Development, 3(1):??, January 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/

html/features/archive/9801/
feigenbaum_main.shtml.

Feigenbaum:1998:CJCb

[Fei98b] Barry Feigenbaum. Code from:

Java class builder. Java Report: The Source for Java Development, 3(1):??, January 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9801/
builder.shtml.

Feigenbaum:1998:FC

[Fei98c] Barry Feigenbaum. Feigenbaum code. Java Report: The Source for Java Development, 3(1):??, January 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9801/
feigenbaumcode.shtml.

Feigenbaum:1998:JCB

[Fei98d] Barry Feigenbaum. Java class builder: An approach to adding a build facility tailored to Java and related languages. Java Report: The Source for Java Development, 3(1):??, January 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9801/
feigenbaum.shtml.

Fitzgerald:1996:PRS

[Fit96] Patrick Fitzgerald. Product review: Symantec’s Café for the Macintosh. Java Report: The Source for Java Development, 1 (3):23–24, July and August 1996.

CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9607/java9607.toc.html.

Flynn:1996:PRM

[Fly96] Jim Flynn. Product review: Microsoft’s Visual J++. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CO-DEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9611/java9611.toc.html.

Flynn:1997:WYN

[Fly97] Jim Flynn. What you need to know about Microsoft’s Authenticode. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

                                        //www.sigs.com/publications/
                                        docs/java/9702/java9702.toc.
                                        html.

Frankel:1998:TIB

[Fra98] David S. Frankel. Technologyindependent business objects — the concept of a meta-model. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660.

Gamma:1996:ADP

[Gam96] Erich Gamma. Applying design patterns in Java. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9611/java9611.toc.html.

Geary:1996:GJE

[Gea96] David Geary. Graphic Java:

Embedding images directly in a class. Java Report: The Source for Java Development, 1 (6):??, November/December 1996. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9611/java9611.toc.html.

Geary:1997:GJAb

[Gea97a] David Geary. Graphic Java: 1.1 AWT makes strides. Java Report: The Source for Java Development, 2(8):??, September 1997. CODEN JREPFI. ISSN 1086-4660.

Geary:1997:GJAa

[Gea97b] David Geary. Graphic Java: Answering frequently asked AWT questions. Java Report: The Source for Java Development, 2 (2):??, February 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9702/java9702.toc.html.

Geary:1997:GJJ

[Gea97c] David Geary. Graphic Java:

JFC’s Swing, part 1: Model/ view/controller. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9711/java9711.toc.html.

Geary:1997:GJM

[Gea97d] David Geary. Graphic Java: Model/view/controller. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9701/java9701.toc.html.

Geary:1997:GJEa

[Gea97e] David Geary. Graphic Java: The 1.1 event handling model. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9703/java9703.toc.
html;   http://www.sigs.com/
publications/docs/java/9703/
9703.geary.html.

Geary:1997:GJEb

[Gea97f] David Geary. Graphic Java: The 1.1 event handling model. Java Report: The Source for Java Development, 2(5):66–??, May 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Geary:1998:SM

[Gea98] David Geary. Swing and multithreading. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9811/
html/from_pages/ftp_col1.shtml.

Gill:1996:ENJ

[Gil96] Philip Gill. Editor’s note: Java buzz. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9611/java9611.toc.html.

Gill:1997:JBa

[Gil97a] Philip Gill. Java buzz. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9702/java9702.toc.html.

Gill:1997:JBb

[Gil97b] Philip Gill. Java buzz. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9703/java9703.toc.html.

Gill:1997:JBc

[Gil97c] Philip Gill. Java buzz. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Gill:1997:CSJ

[Gil97d] Philip J. Gill. Case study: Java in the financial services industry. Java Report: The Source for Java Development, 2(4):??, April 1997.

CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Gill:1997:CRJ

[Gil97e] Philip J. Gill. Conference report: The Java revolution: Year two. Java Report: The Source for Java Development, 2(5):14–??, May 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9705/9705.toc.html.

Gill:1997:CPC

[Gil97f] Philip J. Gill. Corporate profile: Corel aims at Microsoft, hopes to overtake software giant with All-Java Office Suite. Java Report: The Source for Java Development, 2(7):??, July/August 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9707/9707.toc.html.

Gill:1997:CPH

[Gil97g] Philip J. Gill. Corporate profile: HP coming on strong in Java. Java Report: The Source for Java Development, 2(6):16– ??, June 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9706/java9706.toc.
html.

Gill:1997:CPJ

[Gil97h] Philip J. Gill. Corporate profile: JAVA key to Borland resurgence? Java Report: The Source for Java Development, 2(5):18–??, May 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9705/9705.toc.html.

Gill:1997:CPS

[Gil97i] Philip J. Gill. Corporate profile: Sanga seeks enterprise beachhead. Java Report: The Source for Java Development, 2(11):16– ??, December 1997. CODEN JREPFI. ISSN 1086-4660.

Gill:1997:JBd

[Gil97j] Philip J. Gill. Java buzz. Java Report: The Source for Java Development, 2(5):8–??, May 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Gill:1997:JBe

[Gil97k] Philip J. Gill. Java buzz. Java Report: The Source for Java Development, 2(6):6–??, June 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/java9706.toc.html.

Gill:1997:JBf

[Gil97l] Philip J. Gill. Java buzz. Java Report: The Source for Java Development, 2(7):??, July/August 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9707/9707.toc.html.

Gill:1997:JBg

[Gil97m] Philip J. Gill. Java buzz. Java Report: The Source for Java Development, 2(8):8–??, September 1997. CODEN JREPFI. ISSN 1086-4660.

Gill:1997:JBh

[Gil97n] Philip J. Gill. Java buzz. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9711/java9711.toc.html.

Gill:1997:JBi

[Gil97o] Philip J. Gill. Java buzz. Java Report: The Source for Java Development, 2(11):10–??, December 1997. CODEN JREPFI. ISSN 1086-4660.

Gill:1997:PFG

[Gil97p] Philip J. Gill. Product focus: GemStone bets on mid-tier Java. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

                                        sigs.com/publications/docs/
                                        java/9711/java9711.toc.html.

Gill:1998:TTM

[Gil98a] Philip Gill. A three-tiered model for cash management systems. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

                                        archive.javareport.com/9810/
                                        html/from_pages/ftp_realworld.
                                        shtml.

Gill:1998:HCO

[Gil98b] Philip J. Gill. Here come OTMs!; IBM using Java to push CICS into OTM space; Java immunizes Y2K non-compliance?; new SDK targets Java network devices; Java to take lead in Web programming; Sun buys NetDynamics; EDS developing 100% Pure Java business apps; and more. Java Report: The Source for Java Development, 3 (??):??, September 1998. CODEN JREPFI. ISSN 1086-4660.

Gill:1998:IBL

[Gil98c] Philip J. Gill. IBI brews legacy integration beans; Fujitsu has Java-based workflow development platform; Persistence ships EJB app server; EC Cubed’s ecWorkRouter gets 100% Pure Java certified; and more. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9811/
html/from_pages/ftp_javabuzz.
shtml.

Gill:1998:SDJ

[Gil98d] Philip J. Gill. Sun dreams of Jini; Visient offers free lightweight EJB server; new PJava is here!; fastest Java compiler?; IBM, Sun set JavaOS for business release; browser battle nears parity; and more. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9810/
html/from_pages/ftp_javabuzz.
shtml.

Grau:1997:PNa

[Gra97a] Jane Grau. Product news. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

                                        sigs.com/publications/docs/
                                        java/9702/java9702.toc.html.

Grau:1997:PNb

[Gra97b] Jane Grau. Product news. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9703/java9703.toc.html.

Grau:1997:PNc

[Gra97c] Jane Grau. Product news. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9704/9704.toc.html.

Grau:1997:PNd

[Gra97d] Jane Grau. Product news. Java Report: The Source for Java Development, 2(5):74–??, May 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9705/9705.toc.html.

Grau:1997:PNe

[Gra97e] Jane Grau. Product news. Java Report: The Source for Java Development, 2(6):76–??, June 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9706/java9706.toc.html.

Grau:1997:PNf

[Gra97f] Jane Grau. Product news. Java Report: The Source for Java Development, 2(7):??, July/August 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9707/9707.toc.html.

Grau:1997:PNg

[Gra97g] Jane Grau. Product news. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9711/java9711.toc.html.

Grau:1997:PNV

[Gra97h] Jane Grau. Product news: Virtu-Flex Software’s VirtuFlex 2.0; Digitivity Inc.’s CAGE; Visix Software’s Vibe DE; MerzCom Inc.’s MerzScope; Rogue Wave Software’s JDBTools 2.0; RandomNoise Inc.’s Coda; and TV Objects Corp.’s Applet Designer Enterprise. Java Report: The Source for Java Development, 2 (8):73–??, September 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9709/java9709.toc.html.

Houlding:1998:PJB

[Hou98a] Dave Houlding. Putting Java Beans to the test. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660.

Houlding:1998:TCD

[Hou98b] Dave Houlding. Taming CORBA dynamic invocation with Java Beans. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9810/
html/from_pages/ftp_feature.
shtml.

Howard:1997:SJC

[How97] Robert “Rock” Howard. StoreTable: A Java class for simple object persistence. Java Report: The Source for Java Development, 2(5):45–??, May 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Heller:1997:TPC

[HR97] Philip Heller and Simon Roberts. Tips on preparing for the certification exam. Java Report: The Source for Java Development, 2(7):??, July/August 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9707/9707.toc.html.

Hunt:1997:CMU

[Hun97] John Hunt. Constructing modular user interfaces in Java. Java Report: The Source for Java Development, 2(8):25–??, September 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9709/java9709.toc.html.

Jaeschke:1997:JGI

[Jae97] Rex Jaeschke. Java goes international. Java Report: The Source for Java Development, 2 (7):??, July/August 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9707/9707.toc.html.

SIGS:1996:JRS

[Jav96] Java Report: The Source for Java Development, 1996. ISSN 1086-4660. SIGS Books, 71 W 23rd St., New York, NY 10010, USA.

Jepson:1997:DCL

[Jep97] Brian Jepson. Database connectivity: The lure of Java. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9701/java9701.toc.
html;   http://www.sigs.com/
publications/docs/java/9701/
jepson.html.

Jolin:1998:JAA

[Jol98] Art Jolin. Java’s atomic assignment — the key to simpler data access across threads. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660.

Jordan:1997:OJI

[Jor97a] David Jordan. The ODMG Java Interface: An introduction. Java Report: The Source for Java Development, 2(8):45–??, September 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9709/java9709.toc.html.

Jordan:1997:PFS

[Jor97b] David Jordan. Product focus: Simplifying database application development. Java Report: The Source for Java Development, 2 (10):??, November 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9711/java9711.toc.html.

Jordan:1998:AOD

[Jor98a] David Jordan. The 2nd Annual Online Database Derby. Java Report: The Source for Java Development, 3(??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9809/
html/from_pages/.

Jordan:1998:HDP

[Jor98b] David Jordan. A higher degree of portability using ODMG. Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660.

Jovin:1996:DDA

[Jov96] Diana Jovin. Developing database applications in Java: Exploring database-driven Web applications: Java and emerging Internet strategies: reduce cost, improve communication, and capture new revenue. Java Report: The Source for Java Development, 1(3):52–56, July and August 1996. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9607/java9607.toc.html.

Joyce:1996:BRJ

[Joy96]  Michael Joyce.  Book re-

view: Java for C/C++ Programmers. Java Report: The Source for Java Development, 1 (6):??, November/December 1996. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9611/java9611.toc.html.

Juch:1997:SAN

[Juc97] Jay Juch. State of the art:

Network applets behind HTTP proxies. Java Report: The Source for Java Development, 2 (7):??, July/August 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9707/9707.toc.html.

Kalaher:1998:CEU

[Kal98]  Patrick B. Kalaher.  Cod-

ing with the enemy? — understanding how ActiveX integrates with Java on the desktop. Java Report: The Source for Java Development, 3(3):??, March 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9803/
kalaher.shtml.

Kaplan:1997:JMB

[Kap97a] Lowell Kaplan. Java means business: Facilitating enterprise development. Java Report: The Source for Java Development, 2 (10):??, November 1997. CO-DEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9711/java9711.toc.html.

Kaplan:1997:UYW

[Kap97b] Lowell Kaplan. Upgrading your Web site to Java: The first step. Java Report: The Source for Java Development, 2(5):51–??, May 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

                                        sigs.com/publications/docs/
                                        java/9705/9705.toc.html.

Kaplan:1998:JWO

[Kap98a] Lowell Kaplan. Java: Write once, run where??? Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

                                        archive.javareport.com/9810/
                                        html/from_pages/javameansb.
                                        shtml.

Kaplan:1998:PGI

[Kap98b] Lowell Kaplan. Programming with generic interfaces. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

                                        archive.javareport.com/9808/
                                        html/from_pages/ftp_col1.shtml.

Kuznetsov:1997:OPE

[Kuz97] Eugene Kuznetsov. Optimizing performance execution with Java. Java Report: The Source for Java Development, 2(6):49–??, June 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/9706.kuznetsov.html;
http://www.sigs.com/publications/
docs/java/9706/java9706.toc.
html.

Larman:1998:CMW

[Lar98a] Craig Larman. The conceptual model: What’s the object? Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660.

Larman:1998:UCM

[Lar98b] Craig Larman. The use case model: What are the processes? Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660.

Lewis:1998:WMC

[Lew98] Bil Lewis. Writing more complex synchronization in Java. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660.

Major:1997:ENW

[Maj97] Kathleen M. Major. Editorial:

Networking the world. Java Report: The Source for Java Development, 2(5):4–??, May 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Malks:1997:UJW

[Mal97] Danny Malks. Using the Java Web server and the server applet. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9711/java9711.toc.html.

Malarvannan:1998:DYO

[Mal98] Mani Malarvannan. Developing your own protocol handlers in Java. Java Report: The Source for Java Development, 3(2):??, February 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9802/
malarvannan.shtml.

Mangan:1996:JJ

[Man96] Mark Mangan. Java jobs 1.1. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9611/java9611.toc.html.

Martin:1997:JCC

[Mar97] Robert C. Martin. Java and C++: A critical comparison. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9701/java9701.toc.
html;   http://www.sigs.com/
publications/docs/java/9701/
martin.html.

Matsumura:1997:JIB

[Mat97a] Miko Matsumura. Java insider:

Behind Java’s performance boost. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Matsumura:1997:JIH

[Mat97b] Miko Matsumura. Java insider: Hello world. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9702/java9702.toc.html.

McCormick:1996:CPN

[McC96] John McCormick. Corporate profile: Natural intelligence. Java Report: The Source for Java Development, 1(3):20, 22, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

McCormick:1997:DNa

[McC97a] John McCormick. Director’s note.

Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9702/java9702.toc.html.

McCormick:1997:DNb

[McC97b] John McCormick. Director’s note.

Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9703/java9703.toc.html.

McManus:1998:JMM

[McM98] Alex McManus. Java: Memories are made of this. Java Report: The Source for Java Development, 3(??):??, November 1998. CO-DEN JREPFI. ISSN 1086-4660.

Meese:1996:JTI

[Mee96] Philip David Meese. The Java tutor: Improving marquees with double buffering: Rendering images onto graphical areas that aren’t currently visible. Java Report: The Source for Java Development, 1(3):35–39, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Mehler:1997:VNC

[Meh97] Mark Mehler. IS views of network computers. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9703/java9703.toc.
html;   http://www.sigs.com/
publications/docs/java/9703/
9703.mehler.html.

Meyer:1996:TOM

[Mey96] Michel Meyer. True, online multimedia: French experiment with cartoons, but they aren’t fooling around: Getting closer to interactive TV. Java Report: The Source for Java Development, 1 (3):58–59, July and August 1996. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9607/java9607.toc.html.

Meyer:1997:DT

[Mey97] Eric E. Meyer. Development in the trenches. Java Report: The Source for Java Development, 2(4):??, April 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

McGraw:1997:JSH

[MF97] Gary McGraw and Edward Felten. Java security: Hostile applets. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9702/java9702.toc.
html.

Monson-Haefel:1998:EJ

[MH98a] Richard Monson-Haefel. Enterprise JavaBeans. Java Report: The Source for Java Development, 3(3):??, March 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9803/
haefel.shtml.

Monson-Haefel:1998:JND

[MH98b] Richard Monson-Haefel. The Java Naming and Directory Interface (JNDI): A more open and flexible model. Java Report: The Source for Java Development, 3(2):??, February 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

                                        archive.javareport.com/9808/
                                        html/features/archive/9802/
                                        haefel.shtml.

Milletti:1998:MDP

[Mil98] Umberto Milletti. The magic of dynamic page delivery. Java Report: The Source for Java Development, 3(??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

                                        archive.javareport.com/9809/
                                        html/from_pages/.

Misfeldt:1997:IBD

[Mis97]  Trevor Misfeldt.  Incremental

beans development. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9711/java9711.toc.html.

Maine:1997:CT

[ML97] Shane Maine and Mark Lussier. Coffee talk. Java Report: The Source for Java Development, 2(1):??, January 1997. CO-DEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9701/java9701.toc.html.

Monday:1997:JPA

[Mon97] Paul Monday. A Java proxy application: An introduction to URLStreamHandlers. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9711/java9711.toc.html.

Moritz:1997:WWS

[Mor97] Ron Moritz. Why we shouldn’t fear Java. Java Report: The Source for Java Development, 2 (2):??, February 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9702/java9702.toc.html.

Moskowitz:1997:JTG

[Mos97] David Moskowitz. Java training without getting soaked. Java Report: The Source for Java Development, 2(7):??, July/August 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9707/9707.toc.html.

Mannion:1998:PBT

[MP98] Mike Mannion and Roy Phillips.

Prevention is better than A cure. Java Report: The Source for Java Development, 3(??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9809/
html/from_pages/.

Maletic:1997:TYA

[MS97] Greg Maletic and Eric Shapiro.

Turning your applets into JavaBean. Java Report: The Source for Java Development, 2 (8):33–??, September 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9709/java9709.toc.html.

NSTL:1997:PRP

[Nat97] National Software Testing Laboratories. Product review: PowerJ. Java Report: The Source for Java Development, 2(6):12–??, June 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/powerj.html;   http://
www.sigs.com/publications/
docs/java/9706/java9706.toc.
html.

NSTL:1997:PRB

[NST97] NSTL. Product review: Bulletproof’s JDesigner Pro. Java Report: The Source for Java Development, 2(8):??, September 1997. CODEN JREPFI. ISSN 1086-4660.

Oaks:1996:SST

[Oak96a] Scott Oaks. Scott’s solutions: Two techniques for handling events: A few practical considerations. Java Report: The Source for Java Development, 1 (3):80, 78, July and August 1996. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9607/java9607.toc.html.

Oaks:1996:SSU

[Oak96b] Scott Oaks. Scott’s solutions: Using a URLConnection. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9611/java9611.toc.html.

Oaks:1997:SSS

[Oak97a] Scott Oaks. Scott’s solutions: An SMTP-based mail client. Java Report: The Source for Java Development, 2(8):??, September 1997. CODEN JREPFI. ISSN 1086-4660.

Oaks:1997:SSC

[Oak97b] Scott Oaks. Scott’s solutions:

Changing behavior with a factory. Java Report: The Source for Java Development, 2(5):80–??, May 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9705/9705.toc.html.

Oaks:1997:SSD

[Oak97c] Scott Oaks. Scott’s solutions: Dynamic method invocation. Java Report: The Source for Java Development, 2(6):80–??, June 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/java9706.toc.html.

Oaks:1997:SSE

[Oak97d] Scott Oaks. Scott’s solutions:

Enumerating types within Java. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9702/java9702.toc.html.

Oaks:1997:SSH

[Oak97e] Scott Oaks. Scott’s solutions: How packages work in 1.1. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9704/9704.toc.html.

Oaks:1997:SSPa

[Oak97f] Scott Oaks. Scott’s solutions: In praise of layout managers. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9701/java9701.toc.html.

Oaks:1997:SSJ

[Oak97g] Scott Oaks. Scott’s solutions: Is Java really secure? Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9703/java9703.toc.html.

Oaks:1997:SSN

[Oak97h] Scott Oaks. Scott’s solutions: No one answer for every situation. Java Report: The Source for Java Development, 2 (7):??, July/August 1997. CO-DEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9707/9707.toc.html.

Oaks:1997:SSPc

[Oak97i] Scott Oaks. Scott’s solutions:

Programming with threads in Java 1.2. Java Report: The Source for Java Development, 2(11):88– ??, December 1997. CODEN JREPFI. ISSN 1086-4660.

Oaks:1997:SSPb

[Oak97j] Scott Oaks. Scott’s solutions:

Protecting your byte code. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9711/java9711.toc.html.

Oaks:1998:BBN

[Oak98a] Scott Oaks. Back to basics: Network servers in Java. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660.

Oaks:1998:BSH

[Oak98b] Scott Oaks. The burden of synchronization: Hashtable vs. hashmap. Java Report: The Source for Java Development, 3 (??):??, August 1998. CODEN JREPFI. ISSN 1086-4660.

OBrien:1997:MJO

[O’B97] Patrick O’Brien. Making Java objects persistent. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9701/java9701.toc.
html;   http://www.sigs.com/

publications/docs/java/9701/
obrien.html.

Opderbeck:1997:UNG

[Opd97] Holger Opderbeck. Using next generation technology for Java education. Java Report: The Source for Java Development, 2 (7):??, July/August 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9707/9707.toc.html.

Pepper:1997:CPR

[Pep97] Jon Pepper. Corporate profile: RogueWave delivers the tools. Java Report: The Source for Java Development, 2(6):20–??, June 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9706/java9706.toc.html.

Platt:1997:PRN

[Pla97] Steve Platt. Product review: Netiva’s Web database application. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9711/java9711.toc.html.

Reese:1998:STT

[Ree98] George Reese. The secrets of a three-tier distributed architecture. Java Report: The Source for Java Development, 3(4):??, April 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9804/
reese.shtml.

Roth:1998:IEJ

[Rot98] Bill Roth. An introduction to Enterprise Java Beans technology. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660.

Schwartz:1997:BRK

[Sch97] Steven H. Schwartz. Book review:

Kick Ass Java by Tonny Espeset. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9703/java9703.toc.
html.

Scotkin:1996:HWA

[Sco96a] Joel Scotkin. How 1.1 will affect business. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9611/java9611.toc.
html;   http://www.sigs.com/
publications/docs/java/9611/
java9611.scotkin.html.

Scotkin:1996:JMB

[Sco96b] Joel Scotkin. Java means business: When do you use Java? answering the business software development question. Java Report: The Source for Java Development, 1(3):74, 76, 77, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/

docs/java/9607/java9607.toc.
html.

Scott:1996:EVW

[Sco96c] Adrian Scott. Enter the virtual world: Creating client-side VRML with JavaScript. Java Report: The Source for Java Development, 1(3):40–45, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Scott:1996:MJV

[Sco96d] Adrian Scott. The marriage of Java and VRML. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9611/java9611.toc.
html;   http://www.sigs.com/
publications/docs/java/9611/
java9611.scott.html.

Scotkin:1997:JMBa

[Sco97a] Joel Scotkin. Java means business: Assessing the Java security layers. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9702/java9702.toc.
html.

Scotkin:1997:JMBc

[Sco97b] Joel Scotkin. Java means business: The best tool for the job. . .. Java Report: The Source for Java Development, 2(5):61–??, May 1997.

CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Scotkin:1997:JMBb

[Sco97c] Joel Scotkin. Java means business: Visualization: How Java can add real value. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9703/java9703.toc.html.

Seidman:1997:IAN

[Sei97] Gerry Seidman. Internet address:

Networking in Java. Java Report: The Source for Java Development, 2(5):31–??, May 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Seshadri:1997:DJC

[Ses97] Govind Seshadri. Distributed Java computing using RMI. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9711/java9711.toc.html.

Seshadri:1998:CS

[Ses98a] Govind Seshadri. Coffee’s served.

Java Report: The Source for Java Development, 3(1):??, January 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/

html/features/archive/9801/
editorial.shtml.

Seshadri:1998:FFF

[Ses98b] Govind Seshadri. Fighting fire with fire. Java Report: The Source for Java Development, 3(2):??, February 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9802/
editorial.shtml.

Seshadri:1998:EDK

[Ses98c] Govind Seshadri. From the Editor’ desk: King CORBA. Java Report: The Source for Java Development, 3(5):??, May 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9805/
editorial.shtml.

Seshadri:1998:FSD

[Ses98d] Govind Seshadri. The future of software development: Assembly-line computing. Java Report: The Source for Java Development, 3(3):??, March 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9803/
editorial.shtml.

Seshadri:1998:UJJ

[Ses98e] Govind Seshadri. Ultimate Java junkie survey. Java Report: The Source for Java Development, 3(4):??, April 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/

html/features/archive/9804/
junkie.shtml.

Sevareid:1997:JND

[Sev97] Jeremy Sevareid. The JDK 1.1’s new delegation event model. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Shafer:1997:JT

[Sha97a] Dan Shafer. JavaScript tips.

Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9703/java9703.toc.
html;   http://www.sigs.com/
publications/docs/java/9703/
9703.shafer.html.

Shah:1997:SAJ

[Sha97b] Apu Shah. State of the art:

JavaSpaces. Java Report: The Source for Java Development, 2(5):16–??, May 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Smith:1998:DOC

[Smi98] Andrew Smith. Distributed observer chains. Java Report: The Source for Java Development, 3(??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9809/
html/from_pages/.

Sommers:1998:CSP

[Som98] Bret Sommers. Creating signed, privileged applets. Java Report: The Source for Java Development, 3(2):??, February 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

                                        archive.javareport.com/9808/
                                        html/features/archive/9802/
                                        somers.shtml.

Shah:1997:SAH

[SR97] Apu Shah and Steve Renaker. State of the art: How applets are being used… and misused. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9704/9704.toc.html.

Sutter:1998:JDG

[Sut98]  Ryan Sutter.  Java devel-

opers get political: An interview with Rick Ross and Madhu Siddalingaiah of the Java Lobby. Java Report: The Source for Java Development, 3(6):??, June 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

                                        archive.javareport.com/9808/
                                        html/features/archive/9805/
                                        javalobbyinterview.shtml.

Swanson:1997:SA

[Swa97] Scott Swanson. State of the art. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

                                        sigs.com/publications/docs/
                                        java/9702/java9702.toc.html.

Thomas:1997:BRB

[Tho97a] Lowell Thomas. Book review:

Building VRML worlds: How to plan, design, and construct VRML sites. Java Report: The Source for Java Development, 2(6):25–??, June 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/java9706.toc.html.

Thomas:1997:PR

[Tho97b] Lowell Thomas. Product review.

Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9702/java9702.toc.html.

Thomas:1997:JMN

[Tho97c] Lowell E. Thomas. The 1997 Java marketplace: New and notable products. Java Report: The Source for Java Development, 2 (11):51–??, December 1997. CO-DEN JREPFI. ISSN 1086-4660.

Thomas:1997:BRJb

[Tho97d] Lowell E. Thomas. Book review: Java Database Programming with JDBC. Java Report: The Source for Java Development, 2(5):24–??, May 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Thomas:1997:BRJa

[Tho97e] Lowell E. Thomas. Book review:

JavaScript and VBScript by Peter Aiken. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9703/java9703.toc.
html.

Thomas:1997:IAJ

[Tho97f] Lowell E. Thomas. Innovative and affordable Java learning tools. Java Report: The Source for Java Development, 2 (7):??, July/August 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9707/9707.toc.html.

Thomas:1997:PRS

[Tho97g] Lowell E. Thomas. Product review: SuperCede — Java edition. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Thomas:1998:AYN

[Tho98a] Lowell E. Thomas. All you need to build robust component-based software: Imperial Software Technology’s Visaj. Java Report: The Source for Java Development, 3 (??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL

http://archive.javareport.
com/9808/html/products/prod_
rev.shtml.

Thomas:1998:V

[Tho98b] Lowell E. Thomas. Visaj. Java Report: The Source for Java Development, 3(??):??, August 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/products/prod_rev.shtml.

Tillinghast:1996:LI

[Til96a] Tig Tillinghast. A look at the improvements in 1.1. Java Report: The Source for Java Development, 1(6):??, November/December 1996. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9611/java9611.toc.
html;   http://www.sigs.com/
publications/docs/java/9611/
java9611.tillinghast.html.

Tillinghast:1996:MAJ

[Til96b] Tig Tillinghast. Madison Avenue Java: Money between strange bedfellows: One’s specialty should command one’s premium; guidelines for billing services. Java Report: The Source for Java Development, 1(3):62–64, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Tillinghast:1997:MAJc

[Til97a] Tig Tillinghast. Madison Avenue Java: Shockwave vs. ActiveX vs. Java. Java Report: The Source for Java Development, 2(7):??, July/August 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9707/9707.toc.html.

Tillinghast:1997:MAJa

[Til97b] Tig Tillinghast. Madison Avenue Java: The net is falling, the net is falling. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

                                        //www.sigs.com/publications/
                                        docs/java/9701/java9701.toc.
                                        html.

Tillinghast:1997:MAJb

[Til97c] Tig Tillinghast. Madison Avenue Java: The ugly edges around Java. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9704/9704.toc.html.

Tillinghast:1998:LFJ

[Til98a] Tig Tillinghast. Lead, follow, or just get out of the way! Java Report: The Source for Java Development, 3(??):??, November 1998. CODEN JREPFI. ISSN 1086-4660.

Tillinghast:1998:SR

[Til98b] Tig Tillinghast. The solution resolution. Java Report: The Source for Java Development, 3 (??):??, September 1998. CODEN JREPFI. ISSN 1086-4660.

Tropeano:1996:GSO

[Tro96] David Tropeano. Getting started with objects: Class design: Refinements in object-oriented analysis. Java Report: The Source for Java Development, 1(3):31– 34, July and August 1996. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9607/java9607.toc.html.

Vargas:1996:CR

[Var96] Byron Vargas. Conference report: The International Java Developers Conference. Java Report: The Source for Java Development, 1(3):28–30, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Vogel:1998:CT

[Vog98] Andreas Vogel. CORBA talk.

Java Report: The Source for Java Development, 3(6):??, June 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9808/
html/features/archive/9805/
vogel.shtml.

Waters:1997:CTJ

[Wat97a] John K. Waters. Coffee talk:

Java alive with sound: An interview with Thomas Dolby Robertson. Java Report: The Source for Java Development, 2 (7):??, July/August 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9707/9707.toc.html.

Waters:1997:CRL

[Wat97b] John K. Waters. Conference report: Lotus Domino Web Developer’s Conference 1997. Java Report: The Source for Java Development, 2(11):22–??, December 1997. CODEN JREPFI. ISSN 1086-4660.

Waters:1997:CRS

[Wat97c] John K. Waters. Conference report: Seybold Publishing Conference ’97. Java Report: The Source for Java Development, 2 (11):22–??, December 1997. CO-DEN JREPFI. ISSN 1086-4660.

Waters:1997:CP

[Wat97d] John K. Waters. Corporate profile. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

                                        //www.sigs.com/publications/
                                        docs/java/9702/java9702.toc.
                                        html.

Waters:1998:CRL

[Wat98a] John K. Waters. Conference report: Lotus Domino Web Developers’ Conference 1997. Java Report: The Source for Java Development, 3(1):??, January 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

                                        archive.javareport.com/9808/
                                        html/features/archive/9801/
                                        conference.shtml.

Waters:1998:IEI

[Wat98b] John K. Waters. Integrating the enterprise — Inprise’s commitment to simplify distributed applications. Java Report: The Source for Java Development, 3(??):??, October 1998. CODEN JREPFI. ISSN 1086-4660. URL http://

archive.javareport.com/9810/
html/from_pages/.

Werman:1997:DCT

[Wer97a] Aaron Werman. The database connection: A tale of two cities. Java Report: The Source for Java Development, 2(11):30–??, December 1997. CODEN JREPFI. ISSN 1086-4660.

Werman:1997:DCJ

[Wer97b] Aaron Werman. The database connection: J/SQL: Taking the pain out of JDBC. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9703/java9703.toc.html.

Werman:1997:DCY

[Wer97c] Aaron Werman. The database connection: Your first JDBC program. Java Report: The Source for Java Development, 2(6):62–??, June 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/werman.html;   http://
www.sigs.com/publications/
docs/java/9706/java9706.toc.
html.

White:1997:ETC

[Whi97] Eric White. Enumerating types with class. Java Report: The Source for Java Development, 2 (8):51–??, September 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.

                                        com/publications/docs/java/
                                        9709/java9709.toc.html.

Winston:1996:HUM

[WN96] Patrick H. Winston and Sundar Narasimhan. How to use the model-view approach to interface design. Java Report: The Source for Java Development, 1 (6):??, November/December 1996. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9611/java9611.toc.html.

Wong:1996:PPA

[Won96a] Henry Wong. The pragmatic programmer: Another look at layout managers: How to write your very own layout manager. Java Report: The Source for Java Development, 1(3):65, 66, 71–73, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http:

                                        //www.sigs.com/publications/
                                        docs/java/9607/java9607.toc.
                                        html.

Wong:1996:PPH

[Won96b] Henry Wong. The pragmatic programmer: How events are delivered. Java Report: The Source for Java Development, 1 (6):??, November/December 1996. CODEN JREPFI. ISSN 1086-

                                        4660. URL http://www.sigs.
                                        com/publications/docs/java/
                                        9611/java9611.toc.html.

Wong:1997:PPLc

[Won97a] Henry Wong. Pragmatic programmer: A look at image observers. Java Report: The Source for Java Development, 2(4):??, April 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9704/9704.toc.html.

Wong:1997:PPEb

[Won97b] Henry Wong. Pragmatic programmer: Event handling with JavaBeans. Java Report: The Source for Java Development, 2 (8):??, September 1997. CODEN JREPFI. ISSN 1086-4660.

Wong:1997:PPEa

[Won97c] Henry Wong. The pragmatic programmer: Exception handling: more than just an add-on feature. Java Report: The Source for Java Development, 2(1):??, January 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9701/java9701.toc.
html.

Wong:1997:PPI

[Won97d] Henry Wong. Pragmatic programmer: Image observers to media trackers. Java Report: The Source for Java Development, 2(5):56–??, May 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9705/9705.toc.html.

Wong:1997:PPJ

[Won97e] Henry Wong. Pragmatic programmer: JavaBeans 101. Java Report: The Source for Java Development, 2(11):82–??, December 1997. CO-DEN JREPFI. ISSN 1086-4660.

Wong:1997:PPLb

[Won97f] Henry Wong. The pragmatic programmer: Life span of an applet. Java Report: The Source for Java Development, 2(3):??, March 1997. CODEN JREPFI. ISSN 1086-4660. URL http:

//www.sigs.com/publications/
docs/java/9703/java9703.toc.
html;   http://www.sigs.com/
publications/docs/java/9703/
9703.wong.html.

Wong:1997:PPLa

[Won97g] Henry Wong. Pragmatic programmer: Looking at the features of the JDK 1.1. Java Report: The Source for Java Development, 2(2):??, February 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9702/java9702.toc.html.

Wong:1997:PPM

[Won97h] Henry Wong. Pragmatic programmer: Moving on to image producers and image consumers. Java Report: The Source for Java Development, 2(6):28–??, June 1997. CODEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9706/9706.wong.html;   http:
//www.sigs.com/publications/
docs/java/9706/java9706.toc.
html.

Wong:1997:PPP

[Won97i] Henry Wong. The pragmatic programmer: Pitfalls of us-

ing the wait() and notify()

methods. Java Report: The Source for Java Development, 2 (7):??, July/August 1997. CO-DEN JREPFI. ISSN 1086-

4660. URL http://www.sigs.
com/publications/docs/java/
9707/9707.toc.html.

Zhou:1997:UCT

[Zho97] Joe Zhou. Using CAB technology to improve applet performance. Java Report: The Source for Java Development, 2(10):??, November 1997. CODEN JREPFI. ISSN 1086-4660. URL http://www.

sigs.com/publications/docs/
java/9711/java9711.toc.html.

Zino:1998:NJS

[Zin98a] Todd C. Zino. Novera’s jBusiness simplifies deployment to one set of business logic. Java Report: The Source for Java Development, 3 (??):??, October 1998. CODEN JREPFI. ISSN 1086-4660. URL

http://archive.javareport.
com/9810/html/products/prod_
rev.shtml.

Zino:1998:UFJ

[Zin98b] Todd C. Zino. Unique features for Java IDEs, yet quirky throughout ObjectShare’s PARTS for Java Professional. Java Report: The Source for Java Development, 3 (??):??, September 1998. CODEN JREPFI. ISSN 1086-4660. URL

http://archive.javareport.
com/9809/html/products/prod_
rev.shtml.